3.If y=80x-16x2, find the difference ∆y-dy if x=4 and ∆x=-0.2

1
Expert's answer
2021-11-23T09:22:35-0500
"y=80x-16x^2"

"\dfrac{dy}{dx}=y'=80-32x"

"y'(4)=80-32(4)=-48"

"\Delta y=-48(-0.2)=9.6"